\begin{thebibliography}{65}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Afzal et~al.(2023)Afzal, Shardlow, Tuarob, Zaman, Sarwar, Ali,
  Aljohani, Lytras, Nawaz, and Hassan}]{afzal2023generative}
Muhammad~Kashif Afzal, Matthew Shardlow, Suppawong Tuarob, Farooq Zaman, Raheem
  Sarwar, Mohsen Ali, Naif~Radi Aljohani, Miltiades~D. Lytras, Raheel Nawaz,
  and Saeed-Ul Hassan. 2023.
\newblock \href {https://doi.org/10.1007/s12652-023-04584-y} {Generative image
  captioning in urdu using deep learning}.
\newblock \emph{Journal of Ambient Intelligence and Humanized Computing},
  14(6):7719--7731.

\bibitem[{Al-Faruq and Fudholi(2023)}]{10.1063/5.0118155}
Umar Abdul~Aziz Al-Faruq and Dhomas~Hatta Fudholi. 2023.
\newblock \href {https://doi.org/10.1063/5.0118155} {{EfficientNet-Transformer
  for image captioning in Bahasa}}.
\newblock \emph{AIP Conference Proceedings}, 2508(1):020037.

\bibitem[{Al-Malki and Al-Aama(2023)}]{s23083783}
Rasha~Saleh Al-Malki and Arwa~Yousuf Al-Aama. 2023.
\newblock \href {https://doi.org/10.3390/s23083783} {Arabic captioning for
  images of clothing using deep learning}.
\newblock \emph{Sensors}, 23(8).

\bibitem[{Al-muzaini et~al.(2018)Al-muzaini, Al-yahya, and
  Benhidour}]{Al-muzaini2018}
Huda~A. Al-muzaini, Tasniem~N. Al-yahya, and Hafida Benhidour. 2018.
\newblock \href {https://doi.org/10.14569/IJACSA.2018.090610} {Automatic arabic
  image captioning using rnn-lstm-based language model and cnn}.
\newblock \emph{International Journal of Advanced Computer Science and
  Applications}, 9(6).

\bibitem[{B et~al.(2023)B, Gunti, R, and S}]{b_multilingual_2023}
Rohan~Reddy B, Swathi Gunti, Prasanna~Kumar R, and Sridevi S. 2023.
\newblock \href {https://doi.org/10.21203/rs.3.rs-3380598/v1} {Multilingual
  {Image} {Captioning}: {Multimodal} {Framework} for {Bridging} {Visual} and
  {Linguistic} {Realms} in {Tamil} and {Telugu} through {Transformers}}.
\newblock preprint, In Review.

\bibitem[{Barrault et~al.(2018)Barrault, Bougares, Specia, Lala, Elliott, and
  Frank}]{barrault-EtAl:2018:WMT}
Lo\"ic Barrault, Fethi Bougares, Lucia Specia, Chiraag Lala, Desmond Elliott,
  and Stella Frank. 2018.
\newblock \href {http://www.aclweb.org/anthology/W18-6402} {Findings of the
  third shared task on multimodal machine translation}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Shared Task Papers}, pages 304--323, Belgium, Brussels. Association for
  Computational Linguistics.

\bibitem[{Bird(2024)}]{bird-2024-must}
Steven Bird. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.acl-long.797} {Must {NLP} be
  extractive?}
\newblock In \emph{Proceedings of the 62nd Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 14915--14929,
  Bangkok, Thailand. Association for Computational Linguistics.

\bibitem[{Bird and Yibarbuk(2024)}]{bird-yibarbuk-2024-centering}
Steven Bird and Dean Yibarbuk. 2024.
\newblock \href {https://aclanthology.org/2024.eacl-long.50} {Centering the
  speech community}.
\newblock In \emph{Proceedings of the 18th Conference of the European Chapter
  of the Association for Computational Linguistics (Volume 1: Long Papers)},
  pages 826--839, St. Julian{'}s, Malta. Association for Computational
  Linguistics.

\bibitem[{{Bui} et~al.(2023){Bui}, {Hieu Nguyen}, and
  {Nguyen}}]{2023arXiv230504166B}
Doanh~C. {Bui}, Nghia {Hieu Nguyen}, and Khang {Nguyen}. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2305.04166} {{UIT-OpenViIC: A
  Novel Benchmark for Evaluating Image Captioning in Vietnamese}}.
\newblock \emph{arXiv e-prints}, arXiv:2305.04166.

\bibitem[{Choudhury et~al.(2023)Choudhury, Guha, and
  Nandi}]{choudhury-etal-2023-image}
Pankaj Choudhury, Prithwijit Guha, and Sukumar Nandi. 2023.
\newblock \href {https://aclanthology.org/2023.paclic-1.74/} {Image caption
  synthesis for low resource {A}ssamese language using {B}i-{LSTM} with
  bilinear attention}.
\newblock In \emph{Proceedings of the 37th Pacific Asia Conference on Language,
  Information and Computation}, pages 743--752, Hong Kong, China. Association
  for Computational Linguistics.

\bibitem[{Das and Singh(2021)}]{das-singh-2021-image}
Ringki Das and Thoudam~Doren Singh. 2021.
\newblock \href {https://aclanthology.org/2021.icon-main.28} {Image caption
  generation framework for {A}ssamese news using attention mechanism}.
\newblock In \emph{Proceedings of the 18th International Conference on Natural
  Language Processing (ICON)}, pages 231--239, National Institute of Technology
  Silchar, Silchar, India. NLP Association of India (NLPAI).

\bibitem[{Das and Singh(2022)}]{das2022assamese}
Ringki Das and Thoudam~Doren Singh. 2022.
\newblock \href {https://doi.org/10.1007/s11042-022-12042-8} {Assamese news
  image caption generation using attention mechanism}.
\newblock \emph{Multimedia Tools and Applications}, 81(7):10051--10069.

\bibitem[{Dhir et~al.(2019)Dhir, Mishra, Saha, and
  Bhattacharyya}]{dhir2019deep}
Rijul Dhir, Santosh~Kumar Mishra, Sriparna Saha, and Pushpak Bhattacharyya.
  2019.
\newblock A deep attention based framework for image caption generation in
  hindi language.
\newblock \emph{Computaci{\'o}n y Sistemas}, 23(3):693--701.

\bibitem[{Dwarampudi et~al.(2021)Dwarampudi, Rampavan, Sathwik, Reddy, Mishra,
  Singh, and Agrawal}]{9441908}
Jagan Mohan~Reddy Dwarampudi, Damerla Rampavan, Mandavilli Anu~Sandeep Sathwik,
  Kaipu~Nivas Reddy, Vipul~Kumar Mishra, Dilbag Singh, and Apeksha Agrawal.
  2021.
\newblock \href {https://doi.org/10.1109/ICACCS51430.2021.9441908} {Image
  caption generation in telugu}.
\newblock In \emph{2021 7th International Conference on Advanced Computing and
  Communication Systems (ICACCS)}, volume~1, pages 438--443.

\bibitem[{ElJundi et~al.(2020)ElJundi, Dhaybi, Mokadam, Hajj, and
  Asmar}]{eljundi2020resources}
Obeida ElJundi, Mohamad Dhaybi, Kotaiba Mokadam, Hazem~M Hajj, and Daniel~C
  Asmar. 2020.
\newblock Resources and end-to-end neural network models for arabic image
  captioning.
\newblock In \emph{VISIGRAPP (5: VISAPP)}, pages 233--241.

\bibitem[{Elliott et~al.(2017)Elliott, Frank, Barrault, Bougares, and
  Specia}]{elliott-EtAl:2017:WMT}
Desmond Elliott, Stella Frank, Lo\"{i}c Barrault, Fethi Bougares, and Lucia
  Specia. 2017.
\newblock \href {http://www.aclweb.org/anthology/W17-4718} {Findings of the
  second shared task on multimodal machine translation and multilingual image
  description}.
\newblock In \emph{Proceedings of the Second Conference on Machine
  Translation}, pages 215--233, Copenhagen, Denmark. Association for
  Computational Linguistics.

\bibitem[{Elliott et~al.(2016)Elliott, Frank, Sima'an, and
  Specia}]{elliott-EtAl:2016:VL16}
Desmond Elliott, Stella Frank, Khalil Sima'an, and Lucia Specia. 2016.
\newblock \href {http://anthology.aclweb.org/W16-3210} {Multi30k: Multilingual
  english-german image descriptions}.
\newblock In \emph{Proceedings of the 5th Workshop on Vision and Language},
  pages 70--74, Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Funaki and Nakayama(2015)}]{funaki2015image}
Ruka Funaki and Hideki Nakayama. 2015.
\newblock \href {https://doi.org/10.18653/v1/D15-1070} {Image-mediated learning
  for zero-shot cross-lingual document retrieval}.
\newblock In \emph{Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 585--590. Association for Computational
  Linguistics.

\bibitem[{Golech et~al.(2022)Golech, Karacan, Sönmez, and Ayral}]{9988025}
Sina~Berk Golech, Saltuk~Buğra Karacan, Elena~Battini Sönmez, and Hakan
  Ayral. 2022.
\newblock \href {https://doi.org/10.1109/ICECCME55909.2022.9988025} {A complete
  human verified turkish caption dataset for ms coco and performance evaluation
  with well-known image caption models trained against it}.
\newblock In \emph{2022 International Conference on Electrical, Computer,
  Communications and Mechatronics Engineering (ICECCME)}, pages 1--6.

\bibitem[{Grubinger et~al.(2006)Grubinger, Clough, M{\"u}ller, and
  Deselaers}]{grubinger2006iapr}
Michael Grubinger, Paul Clough, Henning M{\"u}ller, and Thomas Deselaers. 2006.
\newblock The iapr tc-12 benchmark: A new evaluation resource for visual
  information systems.
\newblock In \emph{International Workshop OntoImage}, volume~5, page~10.

\bibitem[{Hitschler et~al.(2016)Hitschler, Schamoni, and
  Riezler}]{hitschlerETAL:16}
Julian Hitschler, Shigehiko Schamoni, and Stefan Riezler. 2016.
\newblock Multimodal pivots for image caption translation.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (ACL)}.

\bibitem[{Hodosh et~al.(2013)Hodosh, Young, and
  Hockenmaier}]{hodosh2013framing}
Micah Hodosh, Peter Young, and Julia Hockenmaier. 2013.
\newblock Framing image description as a ranking task: Data, models and
  evaluation metrics.
\newblock \emph{Journal of Artificial Intelligence Research}, 47:853--899.

\bibitem[{Hossain and Hasan(2023)}]{https://doi.org/10.17632/rrv8pbxrxv.1}
Md~Anwar Hossain and Mirza AFM~Rashidul Hasan. 2023.
\newblock \href {https://doi.org/10.17632/RRV8PBXRXV.1} {Banglaview: A bangla
  image captioning dataset}.

\bibitem[{Hu et~al.(2023)Hu, Qiao, Kang, and Shi}]{jimaging9080162}
Wenjin Hu, Lang Qiao, Wendong Kang, and Xinyue Shi. 2023.
\newblock \href {https://doi.org/10.3390/jimaging9080162} {Thangka image
  captioning based on semantic concept prompt and multimodal feature
  optimization}.
\newblock \emph{Journal of Imaging}, 9(8).

\bibitem[{Humaira et~al.(2021)Humaira, Paul, Jim, Ami, and Shah}]{Humaira2021}
Mayeesha Humaira, Shimul Paul, Md~Abidur Rahman~Khan Jim, Amit~Saha Ami, and
  Faisal~Muhammad Shah. 2021.
\newblock \href {https://doi.org/10.14569/IJACSA.2021.0120287} {A hybridized
  deep learning method for bengali image captioning}.
\newblock \emph{International Journal of Advanced Computer Science and
  Applications}, 12(2).

\bibitem[{Ilahi et~al.(2020)Ilahi, Zia, Ehsan, Tabassam, and
  Ahmed}]{ilahi2020efficient}
Inaam Ilahi, Hafiz Muhammad~Abdullah Zia, Ahtazaz Ehsan, Rauf Tabassam, and
  Armaghan Ahmed. 2020.
\newblock \href {https://arxiv.org/abs/2008.01663} {Efficient urdu caption
  generation using attention based lstms}.
\newblock \emph{CoRR}, abs/2008.01663.

\bibitem[{Jaknamon and Marukatat(2022)}]{9960246}
Teetouch Jaknamon and Sanparith Marukatat. 2022.
\newblock \href {https://doi.org/10.1109/iSAI-NLP56921.2022.9960246}
  {Thaitc:thai transformer-based image captioning}.
\newblock In \emph{2022 17th International Joint Symposium on Artificial
  Intelligence and Natural Language Processing (iSAI-NLP)}, pages 1--4.

\bibitem[{Jaruschaimongkol et~al.(2023)Jaruschaimongkol, Satirapiwong,
  Pipatsattayanuwong, Temviriyakul, Sangprasert, and
  Siriborvornratanakul}]{Jaruschaimongkol2023automatic}
Manadda Jaruschaimongkol, Krittin Satirapiwong, Kittipan Pipatsattayanuwong,
  Suwant Temviriyakul, Ratchanat Sangprasert, and Thitirat
  Siriborvornratanakul. 2023.
\newblock \href {https://doi.org/10.1007/s43674-023-00068-w} {Automatic image
  captioning in thai for house defect using a deep learning-based approach}.
\newblock \emph{Advances in Computational Intelligence}, 4(1):1.

\bibitem[{Jayaswal et~al.(2023)Jayaswal, Rani, and
  Kaur}]{jayaswal_comprehensive_2023}
Vishal Jayaswal, Rajneesh Rani, and Jagdeep Kaur. 2023.
\newblock \href {https://doi.org/10.21203/rs.3.rs-3601177/v1} {A
  {Comprehensive} {Survey} on {Image} {Captioning} for {Indian} {Languages}:
  {Techniques}, {Datasets}, and {Challenges}}.
\newblock preprint, In Review.

\bibitem[{Jishan et~al.(2020)Jishan, Mahmud, Azad, Alam, and Khan}]{IJAIN499}
Md~Asifuzzaman Jishan, Khan~Raqib Mahmud, Abul Kalam~Al Azad, Md~Shahabub Alam,
  and Anif~Minhaz Khan. 2020.
\newblock \href {https://doi.org/10.26555/ijain.v6i2.499} {Hybrid deep neural
  network for bangla automated image descriptor}.
\newblock \emph{International Journal of Advances in Intelligent Informatics},
  6(2):109--122.

\bibitem[{Kadaoui et~al.(2025)Kadaoui, Atwany, Al-Ali, Mohamed, Mekky, Tilga,
  Fedorova, Artemova, Aldarmaki, and
  Kementchedjhieva}]{kadaoui2025jeemvisionlanguageunderstandingarabic}
Karima Kadaoui, Hanin Atwany, Hamdan Al-Ali, Abdelrahman Mohamed, Ali Mekky,
  Sergei Tilga, Natalia Fedorova, Ekaterina Artemova, Hanan Aldarmaki, and Yova
  Kementchedjhieva. 2025.
\newblock \href {https://arxiv.org/abs/2503.21910} {Jeem: Vision-language
  understanding in four arabic dialects}.
\newblock \emph{Preprint}, arXiv:2503.21910.

\bibitem[{Kaur et~al.(2021)Kaur, Josan, and Kaur}]{Kaur_Josan_Kaur_2021}
Manleen Kaur, Gurpreet Josan, and Jagroop Kaur. 2021.
\newblock \href
  {https://infocomp.dcc.ufla.br/index.php/infocomp/article/view/1180}
  {Automatic punjabi caption generation for sports images}.
\newblock \emph{INFOCOMP Journal of Computer Science}, 20(1).

\bibitem[{Khan et~al.(2022)Khan, Shifath, and Islam}]{khan-etal-2022-ban}
Mohammad~Faiyaz Khan, S.M. Sadiq-Ur-Rahman Shifath, and Md~Saiful Islam. 2022.
\newblock \href {https://aclanthology.org/2022.lrec-1.740/} {{BAN}-cap: A
  multi-purpose {E}nglish-{B}angla image descriptions dataset}.
\newblock In \emph{Proceedings of the Thirteenth Language Resources and
  Evaluation Conference}, pages 6855--6865, Marseille, France. European
  Language Resources Association.

\bibitem[{Lam et~al.(2020)Lam, Le, Nguyen, and
  Nguyen}]{DBLP:journals/corr/abs-2002-00175}
Quan~Hoang Lam, Quang{-}Duy Le, Kiet~Van Nguyen, and Ngan~Luu{-}Thuy Nguyen.
  2020.
\newblock \href {https://arxiv.org/abs/2002.00175} {Uit-viic: {A} dataset for
  the first evaluation on vietnamese image captioning}.
\newblock \emph{CoRR}, abs/2002.00175.

\bibitem[{Lan et~al.(2017)Lan, Li, and
  Dong}]{DBLP:journals/corr/abs-1708-04390}
Weiyu Lan, Xirong Li, and Jianfeng Dong. 2017.
\newblock \href {https://arxiv.org/abs/1708.04390} {Fluency-guided
  cross-lingual image captioning}.
\newblock \emph{CoRR}, abs/1708.04390.

\bibitem[{Li et~al.(2016)Li, Lan, Dong, and Liu}]{li2016adding}
Xirong Li, Weiyu Lan, Jianfeng Dong, and Hailong Liu. 2016.
\newblock Adding chinese captions to images.
\newblock In \emph{Proceedings of the 2016 ACM International Conference on
  Multimedia Retrieval}, pages 271--275. ACM.

\bibitem[{Li et~al.(2018)Li, Wang, Xu, Lan, Wei, Yang, and
  Xu}]{DBLP:journals/corr/abs-1805-08661}
Xirong Li, Xiaoxu Wang, Chaoxi Xu, Weiyu Lan, Qijie Wei, Gang Yang, and Jieping
  Xu. 2018.
\newblock \href {https://arxiv.org/abs/1805.08661} {{COCO-CN} for cross-lingual
  image tagging, captioning and retrieval}.
\newblock \emph{CoRR}, abs/1805.08661.

\bibitem[{Memiş and Yıldız(2025)}]{MEMIS2025102009}
Abbas Memiş and Serdar Yıldız. 2025.
\newblock \href {https://doi.org/10.1016/j.jestch.2025.102009} {Tiny tr-cap: A
  novel small-scale benchmark dataset for general-purpose image captioning
  tasks}.
\newblock \emph{Engineering Science and Technology, an International Journal},
  64:102009.

\bibitem[{Mishra et~al.(2021{\natexlab{a}})Mishra, Dhir, Saha, and
  Bhattacharyya}]{10.1145/3432246}
Santosh~Kumar Mishra, Rijul Dhir, Sriparna Saha, and Pushpak Bhattacharyya.
  2021{\natexlab{a}}.
\newblock \href {https://doi.org/10.1145/3432246} {A hindi image caption
  generation framework using deep learning}.
\newblock \emph{ACM Trans. Asian Low-Resour. Lang. Inf. Process.}, 20(2).

\bibitem[{Mishra et~al.(2021{\natexlab{b}})Mishra, Dhir, Saha, Bhattacharyya,
  and Singh}]{MISHRA2021107114}
Santosh~Kumar Mishra, Rijul Dhir, Sriparna Saha, Pushpak Bhattacharyya, and
  Amit~Kumar Singh. 2021{\natexlab{b}}.
\newblock \href {https://doi.org/10.1016/j.compeleceng.2021.107114} {Image
  captioning in hindi language using transformer networks}.
\newblock \emph{Computers \& Electrical Engineering}, 92:107114.

\bibitem[{Miyazaki and Shimizu(2016)}]{miyazaki2016cross}
Takashi Miyazaki and Nobuyuki Shimizu. 2016.
\newblock \href {http://www.aclweb.org/anthology/P16-1168} {Cross-lingual image
  caption generation}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1780--1790,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Mobasher et~al.(2022)Mobasher, Zamaninejad, Hashemi, Nobakhtian, and
  Eetemadi}]{mobasher_parsvqa-caps_2022}
Shaghayegh Mobasher, Ghazal Zamaninejad, Maryam Hashemi, Melika Nobakhtian, and
  Sauleh Eetemadi. 2022.
\newblock \href
  {https://www.semanticscholar.org/paper/ParsVQA-Caps%3A-A-Benchmark-for-Visual-Question-and-Mobasher-Zamaninejad/9bfee3dbddb159308f6bc6c2b959e2cb8e53537e?utm_source=alert_email&utm_content=LibraryFolder&utm_campaign=AlertEmails_DAILY&utm_term=LibraryFolder+AuthorCitation&email_index=5-0-5&utm_medium=27682982}
  {{ParsVQA}-{Caps}: {A} {Benchmark} for {Visual} {Question} {Answering} and
  {Image} {Captioning} in {Persian}}.
\newblock In \emph{Proceedings of the Widening Natural Language Processing
  workshop}.

\bibitem[{{Mohamed} et~al.(2023){Mohamed}, {Alwajih}, {Moatez Billah Nagoudi},
  {Alcoba Inciarte}, and {Abdul-Mageed}}]{2023arXiv231108844M}
Abdelrahman {Mohamed}, Fakhraddin {Alwajih}, El~{Moatez Billah Nagoudi},
  Alcides {Alcoba Inciarte}, and Muhammad {Abdul-Mageed}. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2311.08844} {{Violet: A
  Vision-Language Model for Arabic Image Captioning with Gemini Decoder}}.
\newblock \emph{arXiv e-prints}, arXiv:2311.08844.

\bibitem[{Muhammad~Shah et~al.(2021)Muhammad~Shah, Humaira, Jim, Saha~Ami, and
  Paul}]{shah2021bornon}
Faisal Muhammad~Shah, Mayeesha Humaira, Md~Abidur Rahman~Khan Jim, Amit
  Saha~Ami, and Shimul Paul. 2021.
\newblock \href {https://doi.org/10.1007/s42979-021-00975-0} {Bornon: Bengali
  image captioning with transformer-based deep learning approach}.
\newblock \emph{SN Computer Science}, 3(1):90.

\bibitem[{Nath et~al.(2022)Nath, Adhikary, Dadure, Pakray, Manna, and
  Bandyopadhyay}]{nath-etal-2022-image}
Prachurya Nath, Prottay~Kumar Adhikary, Pankaj Dadure, Partha Pakray, Riyanka
  Manna, and Sivaji Bandyopadhyay. 2022.
\newblock \href {https://aclanthology.org/2022.rocling-1.33} {Image caption
  generation for low-resource {A}ssamese language}.
\newblock In \emph{Proceedings of the 34th Conference on Computational
  Linguistics and Speech Processing (ROCLING 2022)}, pages 263--272, Taipei,
  Taiwan. The Association for Computational Linguistics and Chinese Language
  Processing (ACLCLP).

\bibitem[{Nawer et~al.(2023)Nawer, Khan, Alam, Mehedi, and Rasel}]{10099345}
Nafisa Nawer, Md. Shakiful~Islam Khan, MD.~Mustakin Alam, Md~Humaion~Kabir
  Mehedi, and Annajiat~Alim Rasel. 2023.
\newblock \href {https://doi.org/10.1109/CCWC57344.2023.10099345}
  {Transformation of visual information into bangla textual representation}.
\newblock In \emph{2023 IEEE 13th Annual Computing and Communication Workshop
  and Conference (CCWC)}, pages 0436--0441.

\bibitem[{Parida et~al.(2019)Parida, Bojar, and
  Dash}]{hindi-visual-genome:2019}
Shantipriya Parida, Ond{\v{r}}ej Bojar, and Satya~Ranjan Dash. 2019.
\newblock {Hindi Visual Genome: A Dataset for Multimodal English-to-Hindi
  Machine Translation}.
\newblock \emph{Computaci{\'o}n y Sistemas}, 23(4):1499--1505.
\newblock Presented at CICLing 2019, La Rochelle, France.

\bibitem[{Pham et~al.(2024)Pham, Nguyen, Vuong, and Ha}]{pham_ktvic_2024}
Anh-Cuong Pham, Van-Quang Nguyen, Thi-Hong Vuong, and Quang-Thuy Ha. 2024.
\newblock \href {http://arxiv.org/abs/2401.08100} {{KTVIC}: {A} {Vietnamese}
  {Image} {Captioning} {Dataset} on the {Life} {Domain}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2401.08100 [cs].

\bibitem[{Rahman et~al.(2019)Rahman, Mohammed, Mansoor, and
  Momen}]{RAHMAN2019636}
Matiur Rahman, Nabeel Mohammed, Nafees Mansoor, and Sifat Momen. 2019.
\newblock \href {https://doi.org/10.1016/j.procs.2019.06.100} {Chittron: An
  automatic bangla image captioning system}.
\newblock \emph{Procedia Computer Science}, 154:636--642.
\newblock Proceedings of the 9th International Conference of Information and
  Communication Technology [ICICT-2019] Nanning, Guangxi, China January 11-13,
  2019.

\bibitem[{Rajendran et~al.(2016)Rajendran, Khapra, Chandar, and
  Ravindran}]{rajendran2016bridge}
Janarthanan Rajendran, Mitesh~M Khapra, Sarath Chandar, and Balaraman
  Ravindran. 2016.
\newblock Bridge correlational neural networks for multilingual multimodal
  representation learning.
\newblock In \emph{Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}.

\bibitem[{Rathi(2020)}]{9223087}
Ankit Rathi. 2020.
\newblock \href {https://doi.org/10.1109/ICCECE48148.2020.9223087} {Deep
  learning apporach for image captioning in hindi language}.
\newblock In \emph{2020 International Conference on Computer, Electrical \&
  Communication Engineering (ICCECE)}, pages 1--8.

\bibitem[{Saichyshyna et~al.(2023)Saichyshyna, Maksymenko, Turuta, Yerokhin,
  Babii, and Turuta}]{saichyshyna-etal-2023-extension}
Nataliia Saichyshyna, Daniil Maksymenko, Oleksii Turuta, Andriy Yerokhin,
  Andrii Babii, and Olena Turuta. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.unlp-1.7} {Extension
  {M}ulti30{K}: Multimodal dataset for integrated vision and language research
  in {U}krainian}.
\newblock In \emph{Proceedings of the Second Ukrainian Natural Language
  Processing Workshop (UNLP)}, pages 54--61, Dubrovnik, Croatia. Association
  for Computational Linguistics.

\bibitem[{Santos et~al.(2021)Santos, Colombini, and
  Avila}]{santos_pracegover_2021}
Gabriel Oliveira~dos Santos, Esther~Luna Colombini, and Sandra Avila. 2021.
\newblock \href {http://arxiv.org/abs/2103.11474} {\#{PraCegoVer}: {A} {Large}
  {Dataset} for {Image} {Captioning} in {Portuguese}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2103.11474 [cs].

\bibitem[{Thapliyal et~al.(2022)Thapliyal, Pont~Tuset, Chen, and
  Soricut}]{thapliyal-etal-2022-crossmodal}
Ashish~V. Thapliyal, Jordi Pont~Tuset, Xi~Chen, and Radu Soricut. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.emnlp-main.45}
  {Crossmodal-3600: A massively multilingual multimodal evaluation dataset}.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in
  Natural Language Processing}, pages 715--729, Abu Dhabi, United Arab
  Emirates. Association for Computational Linguistics.

\bibitem[{Tien et~al.(2020)Tien, Do, and Nguyen}]{10.1007/978-3-030-63119-2_64}
Ha~Nguyen Tien, Thanh-Ha Do, and Van-Anh Nguyen. 2020.
\newblock Image captioning in vietnamese language based on deep learning
  network.
\newblock In \emph{Advances in Computational Collective Intelligence}, pages
  789--800, Cham. Springer International Publishing.

\bibitem[{Unal et~al.(2016)Unal, Citamak, Yagcioglu, Erdem, Erdem, Cinbis, and
  Cakici}]{unal2016tasviret}
Mesut~Erhan Unal, Begum Citamak, Semih Yagcioglu, Aykut Erdem, Erkut Erdem,
  Nazli~Ikizler Cinbis, and Ruket Cakici. 2016.
\newblock Tasviret: G\"or\"unt\"ulerden otomatik t\"urk\c{c}e a\c{c}{\i}klama
  olu\c{s}turma \.{I}\c{c}in bir denekta\c{c}{\i} veri k\"umesi (tasviret: A
  benchmark dataset for automatic turkish description generation from images).
\newblock In \emph{IEEE Sinyal \.{I}\c{s}leme ve \.{I}leti\c{s}im
  Uygulamalar{\i} Kurultay{\i} (SIU 2016)}.

\bibitem[{van Miltenburg(2019)}]{van2019pragmatic}
Emiel van Miltenburg. 2019.
\newblock \emph{Pragmatic factors in (automatic) image description}.
\newblock Ph.D. thesis, Vrije Universiteit Amsterdam.

\bibitem[{van Miltenburg et~al.(2017)van Miltenburg, Elliott, and
  Vossen}]{vanmiltenburg-elliott-vossen:2017:INLG2017}
Emiel van Miltenburg, Desmond Elliott, and Piek Vossen. 2017.
\newblock Cross-linguistic differences and similarities in image descriptions.
\newblock In \emph{Proceedings of the 10th International Conference on Natural
  Language Generation}, pages 21--30, Santiago de Compostela, Spain.
  Association for Computational Linguistics.

\bibitem[{{van Miltenburg} et~al.(2018){van Miltenburg}, K\'adar, Koolen, and
  Krahmer}]{miltenburg2018DIDEC}
Emiel {van Miltenburg}, \'Akos K\'adar, Ruud Koolen, and Emiel Krahmer. 2018.
\newblock {DIDEC}: The {D}utch {I}mage {D}escription and {E}ye-tracking
  {C}orpus.
\newblock In \emph{Proceedings of COLING 2018, the 27th International
  Conference on Computational Linguistics}.
\newblock Resource available at \url{https://didec.uvt.nl}.

\bibitem[{Viridiano et~al.(2024)Viridiano, Lorenzi, Timponi~Torrent, Matos,
  Pagano, Sathler~Sigiliano, Gamonal, de~Andrade~Abreu, Vicente~Dutra,
  Samagaio, Carvalho, Campos, Azalim, Mazzei, Fonseca~de Oliveira, Luz,
  Padua~Ruiz, Bellei, Pestana, Costa, Rabelo, Silva, Roza, Souza~Mota,
  Oliveira, and Pelegrino~de Freitas}]{viridiano-etal-2024-framed-multi30k}
Marcelo Viridiano, Arthur Lorenzi, Tiago Timponi~Torrent, Ely~E. Matos,
  Adriana~S. Pagano, Nat{\'a}lia Sathler~Sigiliano, Maucha Gamonal, Helen
  de~Andrade~Abreu, L{\'\i}via Vicente~Dutra, Mairon Samagaio, Mariane
  Carvalho, Franciany Campos, Gabrielly Azalim, Bruna Mazzei, Mateus Fonseca~de
  Oliveira, Ana~Carolina Luz, Livia Padua~Ruiz, J{\'u}lia Bellei, Amanda
  Pestana, Josiane Costa, Iasmin Rabelo, Anna~Beatriz Silva, Raquel Roza,
  Mariana Souza~Mota, Igor Oliveira, and M{\'a}rcio~Henrique Pelegrino~de
  Freitas. 2024.
\newblock \href {https://aclanthology.org/2024.lrec-main.656} {Framed
  {M}ulti30{K}: A frame-based multimodal-multilingual dataset}.
\newblock In \emph{Proceedings of the 2024 Joint International Conference on
  Computational Linguistics, Language Resources and Evaluation (LREC-COLING
  2024)}, pages 7438--7449, Torino, Italy. ELRA and ICCL.

\bibitem[{Vishnu~Kumar and Lalithamani(2022)}]{9848810}
V~H Vishnu~Kumar and N~Lalithamani. 2022.
\newblock \href {https://doi.org/10.1109/AIC55036.2022.9848810} {English to
  tamil multi-modal image captioning translation}.
\newblock In \emph{2022 IEEE World Conference on Applied Intelligence and
  Computing (AIC)}, pages 332--338.

\bibitem[{Wijadi et~al.(2024)Wijadi, Ghiffar, and Qomariyah}]{10497459}
Jason~Jeremy Wijadi, Bryn Ghiffar, and Nunung~Nurul Qomariyah. 2024.
\newblock \href {https://doi.org/10.1109/ICoSEIT60086.2024.10497459}
  {Indonesian language image captioning using encoder-decoder with attention
  approach}.
\newblock In \emph{2024 2nd International Conference on Software Engineering
  and Information Technology (ICoSEIT)}, pages 342--347.

\bibitem[{Wu et~al.(2017)Wu, Zheng, Zhao, Li, Yan, Liang, Wang, Zhou, Lin, Fu,
  Wang, and Wang}]{DBLP:journals/corr/abs-1711-06475}
Jiahong Wu, He~Zheng, Bo~Zhao, Yixin Li, Baoming Yan, Rui Liang, Wenjia Wang,
  Shipei Zhou, Guosen Lin, Yanwei Fu, Yizhou Wang, and Yonggang Wang. 2017.
\newblock \href {https://arxiv.org/abs/1711.06475} {{AI} challenger : {A}
  large-scale dataset for going deeper in image understanding}.
\newblock \emph{CoRR}, abs/1711.06475.

\bibitem[{Yoshikawa et~al.(2017)Yoshikawa, Shigeto, and
  Takeuchi}]{yoshikawa2017stair}
Yuya Yoshikawa, Yutaro Shigeto, and Akikazu Takeuchi. 2017.
\newblock Stair captions: Constructing a large-scale japanese image caption
  dataset.
\newblock \emph{arXiv preprint arXiv:1705.00823}.

\bibitem[{Yıldız et~al.(2023)Yıldız, Memiş, and Varlı}]{10286693}
Serdar Yıldız, Abbas Memiş, and Songül Varlı. 2023.
\newblock \href {https://doi.org/10.1109/UBMK59864.2023.10286693} {Automatic
  turkish image captioning: The impact of deep machine translation}.
\newblock In \emph{2023 8th International Conference on Computer Science and
  Engineering (UBMK)}, pages 414--419.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[final]{acl}

\usepackage{times}
\usepackage{latexsym}
\usepackage{enumitem}
\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{graphicx}
\usepackage{booktabs}
\usepackage{amssymb}
\title{Image captioning in different languages}

\author{Emiel van Miltenburg\\
 Tilburg University\\
  \texttt{C.W.J.vanMiltenburg@tilburguniversity.edu}}

\begin{document}
\maketitle
\begin{abstract}
This short position paper provides a manually curated list of non-English image captioning datasets (as of May 2024). Through this list, we can observe the dearth of datasets in different languages: only 23 different languages are represented. With the addition of the Crossmodal-3600 dataset \cite[36 languages]{thapliyal-etal-2022-crossmodal} this number increases somewhat, but still this number is small compared to the $\pm$~500 institutional languages that are out there. This paper closes with some open questions for the field of Vision \& Language.
\end{abstract}

\section{Introduction}
Image captioning datasets have been developed for several different languages, but only a handful (mostly English datasets) have received attention in the NLP community. This research note aims to provide an overview of the different image captioning datasets that have been described in the literature. This list is the result of monitoring Google Scholar alerts for the last eight years using the keywords "MS COCO" and "Flickr30K" (the prototypical English captioning datasets), and manually curating a list of papers that describe image captioning work in languages other than English.

\subsection{Earlier work}
An earlier version of this list appeared in the PhD thesis of \citet{van2019pragmatic}. The list shows the language of the dataset, the source of the images (either an existing dataset such as Flickr8/30K and MS COCO or the name of the newly introduced dataset, whether the captions were translated from English or independently collected, and a reference to the paper where the dataset was introduced.

Next to individual references, I also relied on the survey by \citet{jayaswal_comprehensive_2023} to get an overview of image captioning work for Indian languages (i.e.\ \textit{Assamese, Bengali/Bangla, Hindi, Punjabi, Tamil, Telugu}, and \textit{Urdu}). Although the survey itself does not contain a list of existing datasets, the references provided a useful starting point to add all Indian captioning datasets to the list.

\subsection{Additions}
Given the creation history of this overview, I probably have overlooked some datasets. Let me know if you have developed a captioning dataset in a language other than English.

\begin{table*}
	\small
	\centering
	\begin{tabular}{lllll}
		\toprule
		Language & Source & T & I & Citation\\
		\midrule
        Arabic & JEEM & & $\checkmark$ & \citealt{kadaoui2025jeemvisionlanguageunderstandingarabic}\\
        Arabic & ArabicFashionData & & $\checkmark$ & \citealt{s23083783}\\
        Arabic & MS COCO* & $\checkmark$& &\citealt{eljundi2020resources}\\
        Arabic & MS COCO & $\checkmark$ &  & \citealt{2023arXiv231108844M}\\
        Arabic & MS COCO* &  & $\checkmark$ & \citealt{2023arXiv231108844M}\\
        Arabic & MS COCO* & $\checkmark$ &  & \citealt{Al-muzaini2018}\\
        Arabic & Flickr8K* & $\checkmark$ &  & \citealt{Al-muzaini2018}\\
		Assamese & Assamese-news & & $\checkmark$ & \citealt{das-singh-2021-image,das2022assamese}\\
        Assamese & Flickr30K & $\checkmark$ & & \citealt{nath-etal-2022-image}\\
        Assamese & $\hookrightarrow$ revised & $\checkmark$ & & \citealt{choudhury-etal-2023-image}\\
        Assamese & MS COCO & $\checkmark$ & & \citealt{nath-etal-2022-image}\\
        Assamese & $\hookrightarrow$ revised & $\checkmark$ & & \citealt{choudhury-etal-2023-image}\\
        Bengali & BanglaLekha & & $\checkmark$ & \citealt{RAHMAN2019636}\\
        Bengali & Bornon & & $\checkmark$ & \citealt{shah2021bornon}\\
        Bengali & Flickr8K & $\checkmark$ & & \citealt{shah2021bornon,Humaira2021}\\
        Bengali & Flickr8K & $\checkmark$ & & \citealt{khan-etal-2022-ban}\\
        Bengali & BNLIT & & $\checkmark$ & \citealt{IJAIN499}\\
        Bengali & Flickr30K & $\checkmark$ & & \citealt{10099345}\\
        Bengali & Flickr30K & $\checkmark$ & & \citealt{https://doi.org/10.17632/rrv8pbxrxv.1}\\
        Brazilian Portuguese & Flickr30K & $\checkmark$ & $\checkmark$ & \citealt{viridiano-etal-2024-framed-multi30k}\\
        Chinese & Flickr8K & $\checkmark$ & $\checkmark$ & \citealt{li2016adding} \\
        Chinese & Flickr30K & $\checkmark$ & $\checkmark$ & \citealt{DBLP:journals/corr/abs-1708-04390}\\
		Chinese & MS COCO* & $\checkmark$ & $\checkmark$ & \citealt{DBLP:journals/corr/abs-1805-08661}\\
        Chinese & AI Challenger/ICC & & $\checkmark$ & \citealt{DBLP:journals/corr/abs-1711-06475}\\
        Chinese & Thangka & & $\checkmark$ & \citealt{jimaging9080162}\\
		Czech & Flickr30K & $\checkmark$ & & \citealt{barrault-EtAl:2018:WMT}\\
		Dutch & Flickr30K* & & $\checkmark$ & \citealt{vanmiltenburg-elliott-vossen:2017:INLG2017}\\
		Dutch & MS COCO* & & $\checkmark$ & \citealt{miltenburg2018DIDEC}\\
		French & MS COCO* & $\checkmark$ & & \citealt{rajendran2016bridge}\\
		French & Flickr30K & $\checkmark$ & & \citealt{elliott-EtAl:2017:WMT}\\
		German & MS COCO* & $\checkmark$ & & \citealt{rajendran2016bridge}\\
		German & MS COCO* & $\checkmark$ & & \citealt{hitschlerETAL:16} \\
		German & Flickr30K & $\checkmark$ & $\checkmark$ & \citealt{elliott-EtAl:2016:VL16}\\
		German & IAPR-TC12 & & $\checkmark$ & \citealt{grubinger2006iapr}\\
        Hindi & Flickr8K & $\checkmark$ & & \citealt{9223087}\\
        Hindi & MS COCO & $\checkmark$ & & \citealt{dhir2019deep,10.1145/3432246,MISHRA2021107114}\\	
        Hindi & Visual Genome & $\checkmark$ & & \citealt{hindi-visual-genome:2019}\\
        Indonesian & Flickr8K & $\checkmark$ & & \citealt{10497459}\\
        Indonesian & MS COCO & $\checkmark$ & & \citealt{10.1063/5.0118155}\\
        Japanese & UIUC Pascal & $\checkmark$ & & \citealt{funaki2015image}\\ 
		Japanese & MS COCO* & & $\checkmark$ & \citealt{miyazaki2016cross}\\
		Japanese & MS COCO & & $\checkmark$ & \citealt{yoshikawa2017stair}\\
        Persian & ParsVQA-Caps & & $\checkmark$ & \citealt{mobasher_parsvqa-caps_2022}\\
        Portuguese & \#{}PraCegoVer & & $\checkmark$ & \citealt{santos_pracegover_2021}\\
        Punjabi & Punjabi Tribune & & $\checkmark$ & \citealt{Kaur_Josan_Kaur_2021}\\
		Spanish & IAPR-TC12 & $\checkmark$ & & \citealt{grubinger2006iapr}\\
		\bottomrule
	\end{tabular}
	\caption{Image description datasets available in languages other than English (A-S), with an indication of their source, and whether the descriptions were \textbf{T}ranslated or \textbf{I}ndependently collected. Asterisks indicate that the data is a subset of the original dataset.
	Flickr8K is the predecessor of Flickr30K, see \protect\citealt{hodosh2013framing}. For some recources multiple references are given since it is not clear which is the `main' reference. In the case of Assamese, \citet{choudhury-etal-2023-image} manually corrected translation errors from the datasets provided by \citet{nath-etal-2022-image}, resulting in a revised (but highly similar) dataset.}
	\label{table:otherlanguages}
\end{table*}

\begin{table*}
	\small
	\centering
	\begin{tabular}{lllll}
		\toprule
		Language & Source & T & I & Citation\\
		\midrule
        Tamil & MS COCO & $\checkmark$ & & \citealt{9848810}\\
        Tamil & MS COCO & $\checkmark$ & & \citealt{b_multilingual_2023}\\
        Tamil & Flickr8K & $\checkmark$ & & \citealt{b_multilingual_2023}\\
        Tamil & Flickr30K & $\checkmark$ & & \citealt{b_multilingual_2023}\\
        Telugu & MS COCO & $\checkmark$ & & \citealt{b_multilingual_2023}\\
        Telugu & Flickr8K & $\checkmark$ & & \citealt{b_multilingual_2023}\\
        Telugu & Flickr30K & $\checkmark$ & & \citealt{b_multilingual_2023}\\
        Telugu & Flickr8K & $\checkmark$ & & \citealt{9441908}\\
        Thai & HouseDefects & & $\checkmark$ & \citealt{Jaruschaimongkol2023automatic}\\
        Thai & Thai-Food & & $\checkmark$ & \citealt{9960246}\\
        Thai & Thai-Travel & & $\checkmark$ & \citealt{9960246}\\
        Thai & Flickr30K & $\checkmark$ & & \citealt{9960246}\\
		Turkish & Flickr8K & & $\checkmark$ & \citealt{unal2016tasviret}\\
        Turkish & MS COCO & $\checkmark$ & & \citealt{9988025}\\
        Turkish & MS COCO & $\checkmark$ & & \citealt{10286693}\\
        Turkish & Tiny TR-cap & & $\checkmark$ & \citealt{MEMIS2025102009}\\
        Ukrainian & Flickr30K & $\checkmark$ & & \citealt{saichyshyna-etal-2023-extension}\\
        Urdu & Flickr8K* & $\checkmark$ & & \citealt{ilahi2020efficient}\\
        Urdu & Flickr8K* & $\checkmark$ & & \citealt{afzal2023generative}\\
        Vietnamese & MS COCO* & & $\checkmark$ & \citealt{DBLP:journals/corr/abs-2002-00175}\\
        Vietnamese & KTVIC & & $\checkmark$ & \citealt{pham_ktvic_2024}\\
        Vietnamese & UIT-OpenViIC & & $\checkmark$ & \citealt{2023arXiv230504166B}\\
		\bottomrule
	\end{tabular}
	\caption{Image description datasets available in languages other than English (T-Z), with an indication of their source, and whether the descriptions were \textbf{T}ranslated or \textbf{I}ndependently collected. Asterisks indicate that the data is a subset of the original dataset.
	Flickr8K is the predecessor of Flickr30K, see \protect\citealt{hodosh2013framing}. For some recources multiple references are given since it is not clear which is the `main' reference.}
	\label{table:otherlanguages2}
\end{table*}

\section{The list}
Tables~\ref{table:otherlanguages} and \ref{table:otherlanguages2} provide a list of non-English captioning datasets.\footnote{The Crossmodal-3600 dataset \cite[36 languages]{thapliyal-etal-2022-crossmodal} is not listed in the table because it would result in too much repetition.}$^{,}$\footnote{Some authors opt not to translate the dataset but instead use an English-language captioning model and then translate the results (e.g. \citealt{10.1007/978-3-030-63119-2_64}). These works are also not in the table.} Some observations:
\begin{itemize}
    \item Starting with the most obvious observation: although I made an effort to be inclusive and bookmark all non-English captioning datasets I came across, the table is still not very diverse. Only a handful of languages have researchers actively producing image captioning datasets.

    \item Many datasets have been created through (automatic) translation of English captions, which is relatively cheap compared to independently collected captions. The downside, of course, is that these datasets still capture a Western perspective, even if the dataset has been written in a non-Western language. 

    \item Although Tables~\ref{table:otherlanguages} and \ref{table:otherlanguages2} differentiate between translated and independently collected descriptions, in practice this distinction may be blurred by annotators providing manual corrections of automatic translations, sometimes replacing the translations altogether.

    \item Some datasets are very specific to a particular use case, for example the Arabic Fashion dataset \cite{s23083783}, the Chinese Thangka dataset \cite{jimaging9080162}, and the Thai house defects dataset \cite{Jaruschaimongkol2023automatic}. These may not be useful for general image captioning tasks, but may still help to ground multi-modal language models.

    \item Most of the papers in Table~\ref{table:otherlanguages} have not been published at ACL workshops (or their Computer Vision equivalents), meaning that these are easily missed if you aren't explicitly looking for them.

    \item There are six languages in this list that are not in the Crossmodal-3600 dataset. These are: Assamese, Brazilian Portuguese, Portuguese, Punjabi, Tamil, and Urdu. This brings the total number of languages with available caption data to 36 + 6 = 42.
\end{itemize}

\section{Research questions}
Browsing the list above, many research questions come to mind. For example:

\begin{description}[wide]
    \item[Assessment] How well do existing Vision \& Language models perform on different datasets, such as Assamese-news, BNLIT, or Thai-Food/Travel? And, since any caption task inherently makes (implicit) assumptions about the intended audience, from which perspective should we assess these models?

    \item[Benefiting from multi-lingual data] How can we benefit from multi-lingual datasets? And how do we overcome the challenges brought by the imbalance in available training data?
    
    \item[Inclusion and Community Building] How inclusive are Vision \& Language research and citation practices? Are we doing enough to accommodate and support speakers of underrepresented languages? 
\end{description}

The last point clearly stands out from the previous two, in that it does not only require desk research but also carrying out community work, talking to different stakeholders, and bridging the gap between different cultures. This seems most important to make Vision \& Language models work for everyone, yet this work is not strongly incentivized.

\section{Discussion}
It is often said that there are about 7,000 languages in the world, which may lead us to think that we should also develop NLP technology for all of those languages. But \citet{bird-2024-must} reminds us that beyond the 500 institutional languages, there are 6,500 oral languages that may not require technological solutions such as image captioning.\footnote{See \citet{bird-yibarbuk-2024-centering} for a more in-depth discussion of these numbers.} Instead, the \emph{contact languages} used between `outsiders' and speakers of those oral languages (that overlap with the 500 institutional languages) should be the target language for NLP technology. 

\section{Conclusion}
This note provides an overview of non-English captioning datasets. This is not only useful for those who would like to train a captioning model for the different languages that are listed in the table, but it hopefully also increases the visibility of the authors who have been working to promote image captioning technology for their own languages. It is also clear that much work remains to be done, not only developing datasets and models for different languages and cultures, but also in terms of building a community that embraces cultural diversity.

\bibliography{custom}
\end{document}